The specific heat of a certain type of metal is 0.128 J/(g⋅∘C). What is the final temperature if 305 J of heat is added to 66.5 g of this metal, initially at 20.0 ∘C?
1
Expert's answer
2019-10-17T07:18:12-0400

E = m*Cp*(T2 – T1)

305 = 66.5 x 0.128 x (T2 – 20)

305 + 170.24 = 66.5 x 0.128 x T2

T2 = 475.24/8.512 = 55.83C

55.83C